OpenStudy (anonymous):

find the integral from 0 to 1 of 1/sqrtx

OpenStudy (amistre64):

if this was the integral of 1/2sqrtx it would be easier ...

OpenStudy (amistre64):

\[\int \frac 1{\sqrt x}~dx\] \[\frac22\int \frac 1{\sqrt x}~dx\] \[2\int \frac 1{2\sqrt x}~dx\]

OpenStudy (amistre64):

or just convert sqrt x to x^(-1/2) and use a power rule

OpenStudy (anonymous):

Thank you very much! I was getting caught up in the limits and not paying attention to the math

OpenStudy (amistre64):

youre welcome